### v1

# Pan-LUT: Efficient Pan-sharpening via Learnable Look-Up Tables

Abstract
Recently, deep learning-based pan-sharpening algorithms have achieved notable advancements over traditional methods. However, many deep learning-based approaches incur substantial computational overhead during inference, especially with high-resolution images. This excessive computational demand limits the applicability of these methods in real-world scenarios, particularly in the absence of dedicated computing devices such as GPUs and TPUs. To address these challenges, we propose Pan-LUT, a novel learnable look-up table (LUT) framework for pan-sharpening that strikes a balance between performance and computational efficiency for high-resolution remote sensing images. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Our proposed method contains fewer than 300K parameters and processes a 8K resolution image in under 1 ms using a single NVIDIA GeForce RTX 2080 Ti GPU, demonstrating significantly faster performance compared to other methods. Experiments reveal that Pan-LUT efficiently processes large remote sensing images in a lightweight manner, bridging the gap to real-world applications. Furthermore, our model surpasses SOTA methods in full-resolution scenes under real-world conditions, highlighting its effectiveness and efficiency.
* Equal contributions  $\dagger$ Corresponding author.
1 Introduction
High-resolution multispectral (HRMS) images are widely used in applications such as military operations, environmental monitoring, and mapping. However, due to the limitations of physical sensors, these images are challenging to obtain. Pan-sharpening addresses this issue by fusing high-resolution panchromatic (PAN) images with low-resolution multispectral (LRMS) images, producing high-quality HRMS images through complementary integration [47] [55][61].
Figure 1: Comparisons of performance and computational efficiency. The PSNR value is evaluated on the Worldview-III dataset. Meanwhile, the inference time is measured on images with a resolution of ($2048\times 2048$) by employing a single NVIDIA GeForce RTX 2080 Ti GPU. Our method outperforms traditional approaches by a significant margin of 7 dB, while maintaining a speed comparable to that of traditional methods. Furthermore, our approach even surpasses some DNN-based approaches, outperforming Pan-GAN by more than 0.3 dB while being 145.82 times faster. Note that (4K, 8K) refer to models processing images at resolutions of $4096\times 4096$ and $8192\times 8192$, respectively.
Recently, numerous pan-sharpening methods have been proposed, which can be generally categorized into two main groups: traditional methods and deep learning-based methods. Traditional pan-sharpening methods, such as component substitution (CS) [2] [3] [11], multi-resolution analysis (MRA) [7] [20], and variational optimization (VO) [10] [34], often struggle to restore precise spatial and spectral details in HRMS images. This limitation arises from their heavy reliance on manually designed features and insufficient modeling of prior knowledge. In contrast, deep learning-based pan-sharpening methods have demonstrated exceptional fusion capabilities due to the powerful feature extraction ability of deep neural networks (DNNs). Masi et al. [32] built the Pan-sharpening Neural Network (PNN) model, which first applied CNN to pan-sharpening field, achieving a significant improvement over classical methods. Following this, researchers have explored more complicated and deeper networks to further promote the performance of pan-sharpening [12] [39] [33] [40] [56]. Besides, transformer-based models have garnered attention in pan-sharpening for their remarkable ability in long-range modeling, such as [53] [57] [59]. Nevertheless, many existing approaches suffer from significant computational overhead and memory costs during inference, particularly when dealing with high-resolution images. These excessive demands limit the practical deployment of these methods, especially in environments lacking dedicated computing resources such as GPUs and TPUs. As shown in  Figure 1, very few methods are capable of processing 4K-resolution images under limited computational resources. Even among the deep learning-based methods that can handle 2K-resolution images, their inference speed is often suboptimal. In real-world applications, remote sensing images typically have much higher resolutions, and these methods may encounter significant challenges in operating efficiently under such conditions. Furthermore, simply increasing the depth of a network does not guarantee improved performance, as deeper networks become more difficult to train and may suffer from overfitting due to an excessive number of redundant parameters. To this end, our work aims to overcome these challenges by introducing a more efficient and lightweight pan-sharpening approach that delivers high performance and computational efficiency for high-resolution remote sensing images.
The lookup table (LUT), as a vital constituent in image signal processors (ISPs) [19], owes its significance to several advantageous characteristics, such as runtime efficiency and hardware-friendly nature. They offer a promising data structure for efficiently performing specific transformations by replacing expensive runtime computations with inexpensive array caching and indexing. However, the internal parameters of LUTs are usually preset by experts, making it difficult to optimize them for specific images, which limits their effectiveness.
To overcome the aforementioned challenges, we propose a novel learnable LUT framework, called Pan-LUT, which achieves a good balance between performance and computational efficiency in pan-sharpening. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Our method consists of fewer than 300K parameters and can processes a 8K resolution image in under 1 ms using a single NVIDIA GeForce RTX 2080 Ti GPU. Furthermore, our approach outperforms traditional methods by 7 dB, while maintaining a speed comparable to that of conventional techniques, demonstrating superior speed and efficiency compared to existing methods.
In summary, the contributions are as follows:
  • We present Pan-LUT, a novel learnable LUT framework that does not incorporate any network structure. This framework is designed to achieve a strong balance between performance and computational efficiency in pan-sharpening high-resolution remote sensing images.
  • To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT).
  • To the best of our knowledge, this is the first attempt to introduce LUTs for efficient pan-sharpening. Extensive experiments on different satellite datasets demonstrate the effectiveness and efficiency of Pan-LUT.
Figure 2: The overall framework of our proposed Pan-LUT. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT).
2 Related Works
2.1 Look-Up Table
Lookup Table (LUT) is a widely-used data structure in Image Signal Processors (ISPs) [5] [8] [18], which maps a set of indices to corresponding output values, allowing efficient retrieval without the need for repeated computations. LUTs are particularly useful for functions of multiple variables, as they store precomputed outputs for all possible input combinations. For example, in a 1D LUT, a single input index is mapped to an output value, often using linear interpolation for indices that fall between pre-stored values. More complex LUTs, such as 3D LUTs, use three independent input variables, which may require advanced interpolation methods like trilinear or tetrahedral interpolation. Due to its portability, various LUT based solutions have been proposed for image enhancement [6] [23] [26, 27]. For instance, Zeng et al. [52] and Wang et al. [38] propose image-adaptive 3D LUTs for efficient single-image enhancement. These approaches rely on a network weight predictor to fuse different 3D LUTs, which may pose a limitation on platforms under resource-constrained conditions. Additionally, LUT-based methods have been explored in the area of super-resolution [17] [22] [30] [28]. SRLUT [17] trains a deep super-resolution (SR) network with a restricted receptive field and then cache the output values from the learned SR network in LUTs. However, issues such as performance degradation arise when large patches are cached in LUTs, prompting the development of strategies like MuLUT [22], which introduces multiple LUT variants and a fine-tuning strategy to improve performance. To further enhance the functionality of LUTs, architectures like SPLUT [30] and RCLUT [28] have been proposed. SPLUT processes different image information separately using multiple LUTs, while RCLUT introduces a plugin module to improve LUT-based models with minimal additional computational cost.
2.2 Traditional Pan-sharpening Methods
Traditional fusion techniques encompass component substitution (CS), multi-resolution analysis (MRA), and variational optimization (VO). CS methods, such as IHS [2], Brovey [11], and PCA [3], utilize spatial details from high-resolution panchromatic (PAN) images to replace corresponding details in low-resolution multispectral (LRMS) images, often causing spectral distortion due to the incomplete incorporation of spectral information. MRA techniques, including DWT [20] and ATWT [7], apply multi-resolution decomposition to merge PAN and LRMS images, which helps to reduce spectral distortion. VO methods, such as Bayesian [10] and Total Variation [34], formulate the fusion process as an optimization problem by iteratively minimizing the loss function. While VO methods show promising results, they encounter challenges in optimizing model design and loss functions. Although these approaches have yielded certain improvements, their performance remains constrained by the inadequate modeling, which restricts further advancements in pan-sharpening accuracy and quality.
Table 1: Quantitative comparison across three satellite datasets. The best outcomes are highlighted in red. $\uparrow$ indicates better performance with increasing values, while $\downarrow$ signifies improved performance with decreasing values.
WorldView-II GaoFen2 Worldview-III Inference (ms)
[…]
Brovey [11] 35.8646 0.9216 0.0403 1.8238 37.7974 0.9026 0.0218 1.3720 22.5060 0.5466 0.1159 8.2331 0.20 0.21 0.23
IHS [2] 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616 0.20 0.21 0.23
SFIM [29] 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730 0.26 0.29 0.33
GS [21] 35.6376 0.9176 0.0423 1.8774 37.2260 0.9034 0.0309 1.6736 22.5608 0.5470 0.1217 8.2433 0.68 0.73 0.76
PNN [32] 40.7550 0.9624 0.0259 1.0646 43.1208 0.9704 0.0172 0.8528 29.9418 0.9121 0.0824 3.3206 4.21 12.81 54.59
MSDCNN [50] 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884 16.45 49.82 OOM
Pan-GAN [31] 39.1025 0.9562 0.0303 1.2954 41.4468 0.9661 0.0205 1.0593 28.4959 0.8897 0.0998 3.9067 14.55 40.83 OOM
GPPNN [44] 41.1622 0.9684 0.0244 1.0315 44.2145 0.9815 0.0137 0.7361 30.1785 0.9175 0.0776 3.2593 17.59 43.60 OOM
SFDI [60] 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5462 30.5971 0.9236 0.0741 3.0798 20.26 65.37 OOM
UCGAN [58] 40.0545 0.9553 0.0275 1.1734 42.3634 0.9557 0.0194 0.9480 28.6705 0.8851 0.0990 3.8696 17.26 52.06 OOM
PanFlow [48] 41.8584 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 3.1142 18.27 54.19 OOM
PSCINN [37] 41.8520 0.9703 0.0223 0.9407 47.1100 0.9878 0.0107 0.5612 30.5599 0.9230 0.0748 3.1033 28.31 60.74 OOM
Pan-Mamba [13] 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1551 0.9299 0.0702 2.8942 42.31 87.74 OOM
TA-DiffHQP [41] 42.1255 0.9752 0.0211 0.9023 47.7716 0.9900 0.0101 0.5378 31.3369 0.9302 0.0737 2.6369 481.27 998.58 OOM
Pan-LUT (Ours) 39.8362 0.9555 0.0286 1.1695 42.6559 0.9642 0.0189 0.9044 28.8213 0.8936 0.0935 3.7520 0.27 0.28 0.30
2.3 Deep Learning-based Methods
Deep learning-based methods have emerged as the dominant approach for pan-sharpening in recent years. The PNN [32] model, inspired by SRCNN [9], is the first to introduce CNNs into this domain, surpassing traditional methods. Models like PanNet [49] and MSDCNN [50] further enhance performance by leveraging residual connections and multi-scale convolutions, effectively capturing high-frequency details and supporting a wide range of remote sensing applications. Since then, more complex CNN-based architectures [4] [14] [15] [62] [41] have been proposed in this field to improve pan-sharpening’s mapping ability. Models like GPPNN [44], MMNet [46], and ARFNet [45], which enhance interpretability through deep unfolding techniques and accelerate model convergence. Other approaches, including SFDI [60] and MSDDN [50], utilize Fourier transforms to capture high-frequency features. Transformer-based architectures, such as INN-former [59], Panformer [57] and DRFormer [53] combine CNNs and Transformers to capture both local and global features. Instead of learning a deterministic mapping, generative models such as UCGAN [58], PanFlow [48] and PSCINN [37] generate a distribution of possible outputs for the given inputs. Additionally, some networks [25] [54] [42] [24] focus on improving the generalization capability of CNN-based fusion models by learning generalized features. Despite their promising results, these advanced methods come with high computational costs, limiting their practical applicability.
3 Method
Given the PAN image ($P\in R^{H\times W\times 1}$) and the MS image ($MS\in R^{H/r\times W/r\times C}$), pan-sharpening aims to fuse the complementary information to generate the desirable high spatial resolution MS image $(HRMS\in R^{H\times W\times C})$. Here, $H$ and $W$ denote the height and width of the images, $r$ represents the spatial resolution ratio, with a value of 4, and $C$ denotes the number of spectral bands.
3.1 Framework
The overall framework of our proposed Pan-LUT is illustrated in Figure 2. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Specifically, given the PAN image ($P\in R^{H\times W\times 1}$) and the upsampled MS image ($MS\in R^{H\times W\times C}$), they are first concatenated to form $PM\in R^{H\times W\times(C+1)}$, which is then fed into PGLUT for channel-wise spectral mapping, producing the output $V_{c}\in R^{H\times W\times C}$. SDLUT, using only the PAN image as input, generates local spatial details $V_{s}\in R^{H\times W\times C}$. To further enhance spatial representation, AALUT takes $V_{s}\in R^{H\times W\times C}$ as input to adaptively learn local contexts, resulting in $V_{as}\in R^{H\times W\times C}$. By summing up $V_{c}$, and $V_{as}$, we can obtain the final HRMS result:
$\begin{gathered}V_{c}=PGLUT(PM),\\ V_{s}=SDLUT(P),V_{as}=AALUT(V_{s}),\\ HRMS=V_{c}+V_{as}.\\ \end{gathered}$ (1)
3.2 Spectral Transformation
To preserve the rich spectral information of the MS image, we propose the PAN-guided look-up table (PGLUT), which leverages the PAN image as guidance to finely control the spectral transformation. Specifically, PGLUT is represented as a 5-dimensional matrix containing $N^{5}$ elements, where $N$ denotes the number of bins in each dimension. Each element corresponds to a sampling point, defining a set of indexed input pixels $\{I_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$ and their corresponding output pixels $\{O_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$. Here, $I\in\{pa,r,g,b,nir\}$ represents the pixel values from the PAN and MS images, while $O\in\{R,G,B,NIR\}$ denotes the corresponding cached output pixels.
Since the distribution of LUT elements is discrete in space, the output value cannot be directly retrieved from the LUT. For an input value $\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image. PGLUT first performs a lookup operation to locate the corresponding input pixel in the LUT:
$\begin{gathered}x=\frac{pa_{(w,h)}^{I}}{V_{max}}\cdot N,y=\frac{r_{(w,h)}^{I}}{V_{max}}\cdot N,z=\frac{g_{(w,h)}^{I}}{V_{max}}\cdot N,\\ s=\frac{b_{(w,h)}^{I}}{V_{max}}\cdot N,e=\frac{nir_{(w,h)}^{I}}{V_{max}}\cdot N,\\ \end{gathered}$ (2)
where $V_{max}$ denotes the maximum value (e.g., 255, 1023 or 2047). The coordinates of the sampling points, $L=\{{(i+c,j+c,k+c,m+c,n+c)}\}$, with $c\in\{0,1\}$, can be derived as follows:
$i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,n=\lfloor e\rfloor,$ (3)
[…]
After locating 32 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$O_{(x,y,z,s,e)}=PInterpolation(LUT[L],{\{\mathbf{d}_{l}\}}),$ (4)
where $PInterpolation(\cdot)$ denotes the pentalinear interpolation.
3.3 Spatial Details Transformation
PGLUT is essentially a channel-wise 5D LUT that operates globally, which limits its ability to capture local spatial information. To effectively capture fine-grained spatial details and adaptively learn local contexts, we propose the Spatial Details Lookup Table (SDLUT) and the Adaptive Aggregation Lookup Table (AALUT). As illustrated in Figure 2, given a pixel $p_{(w,h)}$, SDLUT processes this pixel along with its neighboring pixels as input. During the training phase, we employ a rotation ensemble strategy to further expand the receptive field, which can be formulated as:
$\begin{gathered}{V_{1}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}),\\ {V_{2}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w+1,h-1)},p_{(w,h-1)}),\\ {V_{3}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}),\\ {V_{4}}=f(p_{(w,h)},p_{(w+1,h)},p_{(w,h+1)},p_{(w+1,h+1)}).\\ \end{gathered}$ (5)
Subsequently, a learnable AALUT is introduced to adaptively integrates the four outputs. Specially, the AALUT takes the output from SDLUT as input and combines the spatial information to enhance the spatial representation:
${V}_{as}=f_{AALUT}({V_{1}},{V_{2}},{V_{3}},{V_{4}}),$ (6)
where $f_{AVGLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval.
3.4 Loss Function
To achieve satisfying pan-sharpening results, we propose a joint loss for network training. We first utilize the L1 loss:
$\mathcal{L}_{1}=\left\|HRMS-GT\right\|_{1},$ (7)
where $HRMS$ and $GT$ denote the network output and the corresponding ground truth, respectively.
To enhance the stability and robustness of the learned LUTs, we incorporate smoothness regularization $\mathcal{L}_{s}$ and monotonicity regularization $\mathcal{L}_{m}$:
$\mathcal{L}_{s}=\mathcal{L}_{s}^{PG}+\mathcal{L}_{s}^{SD}+\mathcal{L}_{s}^{AA},$ (8)
$\mathcal{L}_{m}=\mathcal{L}_{m}^{PG}+\mathcal{L}_{m}^{SD}+\mathcal{L}_{m}^{AA},$ (9)
where $\mathcal{L}_{s}^{PG}$, $\mathcal{L}_{s}^{SD}$, and $\mathcal{L}_{s}^{AA}$ denote the smoothness regularizations for PGLUT, SDLUT, and AALUT, while $\mathcal{L}_{m}^{PG}$, $\mathcal{L}_{m}^{SD}$, and $\mathcal{L}_{m}^{AA}$ represent the monotonicity regularizations for PGLUT, SDLUT, and AALUT, respectively.
Taking AALUT as an example, the smoothness regularization can be defined as:
$\begin{gathered}\mathcal{L}_{s}^{AA}=\sum_{O\in\{l,o,c,a\}}\sum_{i,j,k,m=0}^{N-1}(\left\|O_{(i+1,j,k,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j+1,k,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j,k+1,m)}-O_{(i,j,k,m)}\right\|^{2}\\ +\left\|O_{(i,j,k,m+1)}-O_{(i,j,k,m)}\right\|^{2}),~{}\end{gathered}$ (10)
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m)}$ is the corresponding output for the defined sampling point $(i,j,k,m)$ in LUT. The definitions of $\mathcal{L}{s}^{PG}$ and $\mathcal{L}{s}^{SD}$ are similar to those in Eq. 10.
The monotonicity regularization in AALUT can be defined as:
$\begin{gathered}\mathcal{L}_{m}^{AA}=\sum_{O\in\{l,o,c,a\}}\sum_{i,j,k,m=0}^{N-1}[g(O_{(i,j,k,m)}-O_{(i+1,j,k,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j+1,k,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j,k+1,m)})\\ +g(O_{(i,j,k,m)}-O_{(i,j,k,m+1)})],~{}\end{gathered}$ (11)
where $g(\cdot)$ denotes the ReLU activation function. Similarly, $\mathcal{L}_{m}^{PG}$ and $\mathcal{L}_{m}^{SD}$ are defined in the same way as in Eq. 11.
The final loss functions are as follows:
$\mathcal{L}=\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{s}+\lambda_{m}\mathcal{L}_{m},$ (12)
where the two constant parameters $\lambda_{s}$ and $\lambda_{m}$ are used to control the effects of the smoothness and monotonicity regularization terms, respectively. In our experiments, we empirically set $\lambda_{s}$ = 0.0001 and $\lambda_{m}$ = 10.
Figure 3: Visual comparison on WorldView-III dataset. The last row visualizes the MSE residues between the pan-sharpening results and the ground truth.
Figure 4: Visual comparison on GaoFen2 dataset. The last row visualizes the MSE residues between the pan-sharpening results and the ground truth.
4 Experiments
In this section, we provide a comprehensive overview of the experimental setup, including the evaluation datasets, performance metrics, and implementation details. We then present both quantitative and qualitative comparisons with other methods. Additionally, we conduct ablation studies to validate the contribution of each component.
4.1 Datasets
Remote sensing datasets from three satellites are used in our experiments, including WorldView-II (WV2), GaoFen2 (GF2) and WorldView-III (WV3). Due to the absence of high-resolution multispectral ground truth images in these datasets, we generate the training set using the Wald protocol tool [36]. Specifically, given the original MS image and its corresponding high-resolution PAN image, they are downsampled by a factor of $r$ to obtain image pairs of MS and PAN, with $r$ set to 4. During training, the original high-resolution MS image is treated as the ground truth, while the MS and PAN images serve as the input image pairs.
Table 2: Evaluation on the real-world full-resolution scenes from WorldView-II dataset. The best values are highlighted by red. The up or down arrow indicates higher or lower metric corresponding to better results.
Metrics Brovey IHS PNN MSDCNN GPPNN SFDI UCGAN PanFlow PSCINN Pan-Mamba TA-DiffHQP Ours
$D_{\lambda}\downarrow$ 0.1026 0.1110 0.1057 0.1063 0.0987 0.1034 0.1042 0.0966 0.0967 0.0966  0.0953 0.0962
$D_{S}\downarrow$ 0.1409 0.1556 0.1446 0.1443 0.1312 0.1305 0.1476 0.1274 0.1271 0.1272  0.1129 0.1253
[…]
Figure 5: Visual comparison on the real full-resolution scenes from the WorldView-II dataset. For a more detailed examination of the results, we zoomed-in view on specific parts of the images.
4.2 Implementation Details
We compare the proposed Pan-LUT against several pan-sharpening methods on reduced-resolution scenes from WV2, WV3, and GF2 datasets. Specifically, we choose four traditional pan-sharpening techniques: Brovey [11], IHS [2], SFIM [29] and GS [21], along with ten deep learning-based approaches: PNN [32], MSDCNN [50], Pan-GAN [31], SFDI [60], UCGAN [58], PanFlow [48], PSCINN [37], Pan-Mamba [13] and TA-DiffHQP [41].
[…]
4.3 Comparison with Other Methods
Evaluation on Reduced-resolution Scene. The quantitative results across three datasets are presented in Table 1, with the best results highlighted in red. Compared to traditional methods, our approach achieves an average PSNR improvement of 4dB, 5dB, and 6dB across the three datasets, while maintaining inference speeds comparable to those of traditional methods. Compared to DNN-based methods, our method exhibits an average PSNR performance decrease of 2 dB, 4 dB, and 2 dB across the three datasets. However, as shown in the table, only PNN is capable of processing 4K-resolution images. Notably, our method is 181.97 times faster than PNN, while the performance gap remains minimal at 0.92 dB, 0.46 dB, and 1.12 dB across the three datasets. It is worth noting that the proposed Pan-LUT does not incorporate any network structure, yet it outperforms some DNN-based methods, such as Pan-GAN and UCGAN, in terms of both performance and inference time. We also provide visual comparisons for the WorldView-III and GaoFen2 datasets, as shown in  Figure 3 and  Figure 4. Evaluation on Full-resolution Scene. To assess the performance and generalization capability of our method on full-resolution scenes under real-world conditions, we first trained Pan-LUT on WorldView-II data and then tested it on unseen full-resolution WorldView-II satellite datasets. The real-world dataset consists of 200 newly collected samples from the WorldView-II satellite for evaluation. The results are presented in Table 2. On reduced-resolution scenes, our method falls short of most DNN-based approaches in terms of performance. However, on full-resolution scenes, it outperforms the majority of them when considering the metrics of $D_{\lambda}$, $D_{S}$, and QNR. This demonstrates its strong generalization ability in real-world situations. Additionally, we provide a visual comparison against both traditional and DNN-based methods, as shown in Figure 5. Evaluation on Reduced-resolution Scene. There are two key limitations of DNN-based methods in practical applications. First, real-world remote sensing satellite images typically possess extremely large sizes, thereby incurring substantial computational burdens. Second, computational resources are often limited, which further restricts the applicability of these methods. Therefore, we conduct comparative experiments to evaluate the computational efficiency of various methods on high-resolution images, including $1024\times 1024$, $2048\times 2048$ and $4096\times 4096$ resolutions. For each method, we record the average inference time on 100 images, with all inference steps conducted on a single NVIDIA GeForce RTX 2080Ti GPU for fair comparison. As shown in Table 1, Pan-LUT efficiently processes images at all resolutions. Compared to DNN-based methods, it achieves significantly faster inference speeds, while maintaining comparable speed to traditional methods. Our method easily meets the real-time processing requirements on GPUs, outperforming all other methods by a substantial margin. Notably, only PNN is capable of handling remote sensing satellite images at the $4096\times 4096$ resolution, highlighting the superior efficiency of our approach.
5 Ablation Study
In this section, we conduct several ablation studies on the WorldView-III dataset to investigate the impact of LUT size and evaluate the effectiveness of each proposed component. Size of Look-Up Tables. As shown in Figure 6, changing the LUT size does not lead to a significant drop in performance. This observation suggests that the effectiveness of our proposed method is not dependent on consuming extensive storage resources to increase the LUT size.
First, we explore the effect of PGLUT size, denoted $N_{P}$. Performance improves with larger $N_{P}$, reaching an optimal point at values $N_{P}$ = 9. Beyond this (from 9 to 17), only a small gain of 0.005 dB is observed, while the parameters increase substantially from 236K to 5M, indicating capacity redundancy. We therefore set $N_{P}$ = 9 as the default to balance performance with storage requirements.
For SDLUT, denoted $N_{S}$. we find that increasing $N_{S}$ from 3 to 9 improves performance, but values above 9 cause a slight performance drop.
Similarly, enlarging AALUT size $N_{A}$ yields only minor gains but substantially increases parameters, especially beyond $N_{A}$ = 9. We set $N_{A}$ = 9 to balance performance with computational efficiency. Effectiveness of Each LUT. We further conduct ablation studies to verify the effectiveness of each LUT. Specifically, we examine the following four settings: (i) only PGLUT, (ii) only SDLUT, (iii) PGLUT and SDLUT, and (iv) SDLUT and AALUT. For (i), the output of PGLUT is directly used as the result. For (ii), the outputs of SDLUT are averaged and then added to the MS image. For (iii), the averaged SDLUT outputs are added to the PGLUT output to generate the result. For (iv), AALUT takes the output of SDLUT as input and performs adaptive aggregation to produce the final result. Results are listed in Table 3. Our observations are as follows: 1) Comparing (i) and (iii), SDLUT effectively captures fine-grained spatial details from the PAN image, thereby enhancing overall performance. 2) Comparing (ii) and (iii), PGLUT provides finer control over spectral transformation, resulting in improved performance. 3) Comparing (ii) and (iv), AALUT demonstrates adaptive aggregation capabilities, effectively learning local contexts and clearly outperforming simple averaging.
Figure 6: Ablation studies on different sizes of the PGLUT, SDLUT and AALUT on the WorldView-III dataset.
Table 3: Ablation study of PGLUT, SDLUT and AALUT on the WorldView-III dataset.
[…]
(i) 23.2254 0.5602 0.1071 7.5064
(ii) 25.9433 0.8132 0.1495 5.5512
(iii) 28.1315 0.8695 0.1183 4.1885
(iv) 26.5035 0.8297 0.1251 5.1258
$\star$ Pan-LUT (Ours) 28.8213 0.8936 0.0935 3.7520
6 Conclusion
In this paper, we propose a novel learnable LUT framework, called Pan-LUT, which strikes an optimal balance between performance and computational efficiency for high-resolution remote sensing images in pan-sharpening. To effectively handle high-resolution images under resource-constrained conditions, we design PGLUT to finely control spectral transformations and facilitate channel-wise spectral mapping. Additionally, we introduce SDLUT and AALUT to capture fine-grained spatial details and adaptively learn local contexts. Extensive experiments on various satellite datasets demonstrate the effectiveness and efficiency of Pan-LUT.