### v2

# Pan-LUT: Efficient Pan-sharpening via Learnable Look-Up Tables

Abstract
Recently, deep learning-based pan-sharpening algorithms have achieved notable advancements over traditional methods. However, deep learning-based methods incur substantial computational overhead during inference, especially with large images. This excessive computational demand limits the applicability of these methods in real-world scenarios, particularly in the absence of dedicated computing devices such as GPUs and TPUs. To address these challenges, we propose Pan-LUT, a novel learnable look-up table (LUT) framework for pan-sharpening that strikes a balance between performance and computational efficiency for large remote sensing images. Our method makes it possible to process $\times15K$ remote sensing images on a 24GB GPU. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details, we introduce the spatial details look-up table (SDLUT). Furthermore, to adaptively aggregate channel information for generating high-resolution multispectral images, we design an adaptive output look-up table (AOLUT). Our model contains fewer than 700K parameters and processes a $\times9K$ image in under 1 ms using one RTX 2080 Ti GPU, demonstrating significantly faster performance compared to other methods. Experiments reveal that Pan-LUT efficiently processes large remote sensing images in a lightweight manner, bridging the gap to real-world applications. Furthermore, our model surpasses SOTA methods in full-resolution scenes under real-world conditions, highlighting its effectiveness and efficiency. The source code is available at https://github.com/CZhongnan/Pan-LUT.
1 Introduction
High-resolution multispectral (HRMS) images are widely used in applications such as military operations, environmental monitoring, and mapping. However, due to the limitations of physical sensors, these images are challenging to obtain. Pan-sharpening addresses this issue by fusing high-resolution panchromatic (PAN) images with low-resolution multispectral (LRMS) images, producing high-quality HRMS images through complementary integration zheng2023deep zhou2023pan wang2023learning. Recently, numerous pan-sharpening methods have been proposed, which can be generally categorized into two main groups: traditional methods and deep learning-based methods. Traditional methods, such as component substitution (CS) ihs pca brovey, multi-resolution analysis (MRA) ATWT DWT, and variational optimization (VO) Bayesian TV, often struggle to restore precise spatial or spectral details in HRMS images. In contrast, deep learning-based pan-sharpening methods have demonstrated exceptional fusion capabilities due to the powerful feature extraction ability of deep neural networks (DNNs) lin2024pair lin2024difftv lin2025jarvisir lin2024aglldiff. Masi et al. pnn built the Pan-sharpening Neural Network (PNN) model, which first applied CNN to pan-sharpening field, achieving a significant improvement over traditional methods.
Figure 1: Comparisons of computational efficiency. Our method can process $\times9K$ and $\times15K$ images on GPUs with 11GB and 24GB memory, respectively. Meanwhile, we observe that (a) DNN-based methods are highly sensitive to the image size, and (b) in the absence of a GPU, they require a considerable amount of time to process images. In the CPU inference time experiments, all methods were conducted on a workstation equipped with an Intel(R) Xeon(R) Gold 6226R CPU.
Following this, researchers have explored more complicated and deeper networks to further promote the performance of pan-sharpening py_2023_ICCV lin2023domain meng2024progressive zhong2025ssdiff. However, they overlook a critical practical issue: the need for real-time processing of large remote sensing images in real-world applications. As illustrated in Figure 1, we observe two major limitations in DNN-based approaches: (1) they are highly sensitive to the size of the input images, and (2) they rely heavily on dedicated computing devices such as GPUs and TPUs. Increasing GPU memory does not significantly improve the image size these methods can handle and several of these methods demand a significant amount of time to process images in CPU-only environments. And we will demonstrate in the experiments sections that, even with GPU acceleration, most methods still fail to process large remote sensing images in real time. In practical applications, remote sensing images typically exhibit even higher resolutions, posing additional challenges to existing methods in terms of efficiency and scalability. Moreover, simply increasing network depth does not necessarily lead to better performance, as deeper models are harder to train and often suffer from overfitting due to redundant parameters.
To overcome the aforementioned challenges, we propose a novel learnable Look-Up Table (LUT) framework, called Pan-LUT, which achieves a good balance between performance and computational efficiency in pan-sharpening. Specifically, we replace complex DNN operations with learnable LUTs to enable lightweight deployment in practical applications. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details, we introduce the spatial details look-up table (SDLUT). To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT). The Pan-LUT consists of fewer than 700K parameters and can processes $\times9K$ images in under 1 ms using a single RTX 2080 Ti GPU. Furthermore, our approach outperforms traditional methods by 7 dB, while maintaining a speed comparable to that of conventional techniques, demonstrating superior speed and efficiency compared to existing methods. Our contributions can be summarized as follows:
  • We present Pan-LUT, a novel learnable LUT framework that does not incorporate any network structure. This framework is designed to achieve a strong balance between performance and computational efficiency in pan-sharpening high-resolution remote sensing images. Our method makes it possible to process $\times15K$ remote sensing images on one 24GB GPU.
  • To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT). To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT).
  • To the best of our knowledge, this is the first attempt to introduce LUTs for efficient pan-sharpening. Extensive experiments on different satellite datasets demonstrate the effectiveness and efficiency of Pan-LUT.
2 Related Work
2.1 Look-Up Table
Look-up Tables (LUT) are particularly useful for functions of multiple variables, as they store precomputed outputs for all possible input combinations. For example, in a 1D LUT, a single input index is mapped to an output value, often using linear interpolation for indices that fall between pre-stored values. More complex LUTs, such as 3D LUTs, use three independent input variables, which may require advanced interpolation methods like trilinear or tetrahedral interpolation. Due to its portability, various LUT based solutions have been proposed for image enhancement conde2024nilut li2023fastllve lin2024unsupervised salut 3dlut. For instance, Zeng et al. 3dlut and Wang et al. salut propose image-adaptive 3D LUTs for efficient single-image enhancement. These approaches rely on a network weight predictor to fuse different 3D LUTs, which may pose a limitation on platforms under resource-constrained conditions. Additionally, LUT-based methods have been explored in the area of super-resolution srlut mulut splut rclut. SRLUT srlut trains a deep super-resolution (SR) network with a restricted receptive field and then cache the output values from the learned SR network in LUTs. However, issues such as performance degradation arise when large patches are cached in LUTs, prompting the development of strategies like MuLUT mulut, which introduces multiple LUT variants and a fine-tuning strategy to improve performance. To further enhance the functionality of LUTs, architectures like SPLUT splut and RCLUT rclut have been proposed. SPLUT processes different image information separately using multiple LUTs, while RCLUT introduces a plugin module to improve LUT-based models with minimal additional computational cost.
Figure 2: The overall framework of our proposed Pan-LUT. PGLUT is a spectral transformation LUT designed to extract spectral information, SDLUT is a spatial detail transformation LUT for capturing texture features, and AOLUT is an adaptive output LUT used to aggregate channel information.
2.2 Traditional Pan-sharpening Methods
Traditional fusion techniques encompass component substitution (CS), multi-resolution analysis (MRA), and variational optimization (VO). CS methods, such as IHS ihs, Brovey brovey, and PCA pca, utilize spatial details from high-resolution panchromatic (PAN) images to replace corresponding details in low-resolution multispectral (LRMS) images, which can lead to spectral distortion due to the incomplete incorporation of spectral information. MRA techniques, including DWT DWT and ATWT ATWT, apply multi-resolution decomposition to merge PAN and LRMS images, which enables better preservation of spectral information and reduces spectral distortion. VO methods, such as Bayesian Bayesian and Total Variation TV, formulate the fusion process as an optimization problem by iteratively minimizing the loss function. While VO methods show promising results, they encounter challenges in optimizing model design and loss functions. Although these approaches have yielded certain improvements, their performance remains constrained by the inadequate modeling, which restricts further advancements in pan-sharpening accuracy and quality.
2.3 Deep Learning-based Methods
Deep learning-based methods have emerged as the dominant approach for pan-sharpening in recent years wang2024cross he2024frequency wang2025towards. The PNN pnn model, inspired by SRCNN srcnn, is the first to introduce CNNs into this domain, surpassing traditional methods. Models like PanNet pannet and MSDCNN msdcnn further enhance performance by leveraging residual connections and multi-scale convolutions, effectively capturing high-frequency details and supporting a wide range of remote sensing applications. Since then, more complex CNN-based architectures chen2024novel hou2023bidomain zhu2023probability have been proposed in this field to improve the mapping ability of pan-sharpening. Models like GPPNN gppnn, MMNet mmnet, and ARFNet arfnet, which enhance interpretability through deep unfolding techniques and accelerate model convergence. Spatial adaptive convolution methods, such as LAGConv jin2022lagconv and CANConv duan2024canconv, can adaptively generate different convolution ker nel parameters based on various spatial locations, enabling them to accommodate different spatial regions. Other approaches, including SFDI sfiin and MSDDN msdcnn, utilize Fourier transforms to capture high-frequency features. Transformer-based architectures, such as INN-former inn, Panformer zhou2022panformer and DRFormer zhang2023drformer combine CNNs and Transformers to capture both local and global features. Instead of learning a deterministic mapping, generative models such as UCGAN ucgan, PanFlow  yang2023panflownet and PSCINN wang2024panpsci generate a distribution of possible outputs for the given inputs. In addition, several studies have investigated lightweight pan-sharpening methods, including SpanConv chen2022spanconv and LGPConv zhao2023lgpconv. However, we show that current lightweight methods are only lightweight in the context of low-resolution images. Despite their promising results, these advanced methods come with high computational costs, limiting their practical applicability.
3 Method
Given the PAN image ($P\in R^{H\times W\times 1}$) and the MS image ($MS\in R^{H/r\times W/r\times C}$), pan-sharpening aims to fuse the complementary information to generate the desirable high spatial resolution MS image $(HRMS\in R^{H\times W\times C})$. Here, $H$ and $W$ denote the height and width of the images, $r$ represents the spatial resolution ratio, with a value of 4, and $C$ denotes the number of spectral bands.
3.1 Framework
The overall framework of our proposed Pan-LUT is illustrated in Figure 2, which consists of three specifically designed LUTs. (1) To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping, which incorporates a PAN-guided indexing strategy and a pentalinear interpolation technique. (2) To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT), which incorporates a rotation-enhanced indexing strategy and a quadrilinear interpolation technique. (3) To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT), which incorporates a PAN-guided indexing strategy and a pentalinear interpolation technique. Specifically, given the PAN image ($P\in\mathbb{R}^{H\times W\times 1}$) and the upsampled MS image ($MS\in\mathbb{R}^{H\times W\times C}$), they are concatenated into $PM\in\mathbb{R}^{H\times W\times(C+1)}$ and passed through the PGLUT for channel-wise spectral mapping, producing the output $V_{pg}\in\mathbb{R}^{H\times W\times(C+1)}$. SDLUT then takes $V_{pg}$ as input to generate local spatial details, yielding $V_{sd}\in\mathbb{R}^{H\times W\times(C+1)}$. Finally, AOLUT takes $V_{sd}$ as input to generate the final $HRMS\in R^{H\times W\times C}$ result:
[…]
3.2 Spectral Transformation
To preserve the rich spectral information of the MS image, we propose the PAN-guided look-up table (PGLUT), which leverages the PAN image as guidance to finely control the spectral transformation. Specifically, PGLUT is represented as a 5-dimensional matrix containing $N^{5}$ elements, where $N$ denotes the number of bins per dimension. Each element corresponds to a sampling point, defining a set of indexed input pixels $\{I_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$ and their corresponding output pixels $\{O_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$. Here, $I\in\{pa,r,g,b,nir\}$ represents the pixel values from the PAN and MS images, while $O\in\{R,G,B,NIR\}$ denotes the corresponding cached output pixels. Since the LUT elements are discretely distributed in space, the output value cannot be directly retrieved from the LUT. For an input value $\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image.
PAN-guided indexing strategy. As illustrated in Figure 2, we introduce an indexing strategy for precise spectral transformation, referred to as the PAN-guided indexing strategy. In a MS image, pixels from different spatial locations may have identical values (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j}$, where $i\neq j$). The LUT maps these identical inputs to the same output. The PAN-guided indexing strategy provides a more flexible indexing mechanism for the LUT. Specifically, it additionally considers the pixels at corresponding spatial positions in the PAN image (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j},pa_{i}\neq pa_{j}$, where $i\neq j$), thereby achieving finer-grained mapping. Specifically, PGLUT first performs a lookup operation to locate the corresponding input pixel in the LUT:
[…]
where $V_{max}$ denotes the maximum value (e.g., 255, 1023 or 2047). The coordinates of the sampling points, $L=\{{(i+c,j+c,k+c,m+c,n+c)}\}$, with $c\in\{0,1\}$, can be derived as follows:
$i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,n=\lfloor e\rfloor,$ (3)
where $\lfloor\cdot\rfloor$ denotes the floor function. $\{\mathbf{d}_{l}\}_{l=x,y,z,s,e}$ represents the offset of the input index $(x,y,z,s,e)$ relative to the defined sampling point $(i,j,k,m,n)$, e.g., $\mathbf{d}_{x}={x}-{i}$.
Pentalinear Interpolation. After locating 32 adjacent points, an appropriate interpolation technique is applied to these sampled values to generate the output:
$O_{(x,y,z,s,e)}=PInterpolation(LUT[L],{\{\mathbf{d}_{l}\}}),$ (4)
where $PInterpolation(\cdot)$ denotes the pentalinear interpolation. More details about PGLUT and the pentalinear interpolation can be found in Section A.2.
3.3 Spatial Details Transformation
PGLUT is essentially a channel-wise 5D LUT that operates globally, which limits its ability to capture local spatial information. To effectively capture fine-grained spatial details and adaptively learn local contexts, we propose the Spatial Details Lookup Table (SDLUT).
Rotation-indexing strategy. As illustrated in Figure 2, given a pixel $p_{(w,h)}$, SDLUT processes this pixel along with its neighboring pixels as input. During the training phase, we employ a Rotation-indexing strategy to further expand the receptive field, which can be formulated as:
[…]
where $f_{SDLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval. More details about SDLUT and the quadrilinear interpolation can be found in Section A.3.
Table 1: Quantitative comparison across three satellite datasets. The best outcomes are highlighted in red. $\uparrow$ indicates better performance with increasing values, while $\downarrow$ signifies improved performance with decreasing values.
WorldView-II GaoFen2 Worldview-III Inference (ms)
[…]
Brovey 35.8646 0.9216 0.0403 1.8238 37.7974 0.9026 0.0218 1.3720 22.5060 0.5466 0.1159 8.2331 - 0.28 0.33
IHS 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616 - 0.23 0.26
SFIM 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730 - 0.32 0.47
GS 35.6376 0.9176 0.0423 1.8774 37.2260 0.9034 0.0309 1.6736 22.5608 0.5470 0.1217 8.2433 - 0.75 0.87
PNN 40.7550 0.9624 0.0259 1.0646 43.1208 0.9704 0.0172 0.8528 29.9418 0.9121 0.0824 3.3206 0.0689 12.81 54.59
PanNet 40.8176 0.9626 0.0257 1.0557 43.0659 0.9685 0.0178 0.8577 29.6840 0.9072 0.0851 3.4263 0.0688 29.52 OOM
MSDCNN 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884 0.2390 49.82 OOM
Pan-GAN 39.1025 0.9562 0.0303 1.2954 41.4468 0.9661 0.0205 1.0593 28.4959 0.8897 0.0998 3.9067 0.0915 40.83 OOM
GPPNN 41.1622 0.9684 0.0244 1.0315 44.2145 0.9815 0.0137 0.7361 30.1785 0.9175 0.0776 3.2593 0.1198 43.60 OOM
SFDI 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5462 30.5971 0.9236 0.0741 3.0798 0.0871 65.37 OOM
UCGAN 40.0545 0.9553 0.0275 1.1734 42.3634 0.9557 0.0194 0.9480 28.6705 0.8851 0.0990 3.8696 0.2109 52.06 OOM
PanFlow 41.8584 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 3.1142 0.0873 54.19 OOM
PSCINN 41.8520 0.9703 0.0223 0.9407 47.1100 0.9878 0.0107 0.5612 30.5599 0.9230 0.0748 3.1033 3.3209 60.74 OOM
Pan-Mamba 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1551 0.9299 0.0702 2.8942 0.1827 87.74 OOM
TA-DiffHQP 42.1255 0.9752 0.0211 0.9023 47.7716 0.9900 0.0101 0.5378 31.3369 0.9302 0.0737 2.6369 2.6000 998.58 OOM
Pan-LUT (Ours) 40.8555 0.9633 0.0254 1.0339 43.7466 0.9726 0.0169 0.8027 29.7376 0.9106 0.0815 3.3934 0.6626 0.38 0.54
3.4 Adaptive Output
For the feature channel pixels from the SDLUT $(V_{(w,h)}^{1},V_{(w,h)}^{2},V_{(w,h)}^{3},V_{(w,h)}^{4},V_{(w,h)}^{5})$, AOLUT adaptively aggregates channel information to generate high-resolution multispectral channel pixels $\{R_{(w,h)}^{O},G_{(w,h)}^{O},B_{(w,h)}^{O},NIR_{(w,h)}^{O}\}$. Pixel-level Transformation can be formulated as:
[…]
where $f_{AOLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval. More details about AOLUT and the pentalinear interpolation can be found in Section A.4.
3.5 Loss Function
To achieve satisfying pan-sharpening results, we propose a joint loss for network training. Suppose the batch size is $T$. We first utilize the MSE loss:
$\mathcal{L}_{mse}=\frac{1}{T}\sum_{t=1}^{T}\|HRMS_{t}-GT_{t}\|^{2},$ (7)
where $HRMS$ and $GT$ denote the network output and the corresponding ground truth, respectively.
To enhance the stability and robustness of the learned LUTs, we incorporate smoothness regularization $\mathcal{L}_{s}$ and monotonicity regularization $\mathcal{L}_{m}$:
$\begin{gathered}\mathcal{L}_{s}=\mathcal{L}_{s}^{PG}+\mathcal{L}_{s}^{SD}+\mathcal{L}_{s}^{AO},\mathcal{L}_{m}=\mathcal{L}_{m}^{PG}+\mathcal{L}_{m}^{SD}+\mathcal{L}_{m}^{AO},\end{gathered}$ (8)
where $\mathcal{L}_{s}^{PG}$, $\mathcal{L}_{s}^{SD}$, and $\mathcal{L}_{s}^{AO}$ denote the smoothness regularizations for PGLUT, SDLUT, and AOLUT, while $\mathcal{L}_{m}^{PG}$, $\mathcal{L}_{m}^{SD}$, and $\mathcal{L}_{m}^{AO}$ represent the monotonicity regularizations for PGLUT, SDLUT, and AOLUT, respectively.
Taking SDLUT as an example, the smoothness regularization can be defined as:
[…]
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m)}$ is the corresponding output for the defined sampling point $(i,j,k,m)$ in LUT. The definitions of $\mathcal{L}{s}^{PG}$ and $\mathcal{L}{s}^{AO}$ are similar to those in Equation 36.
The monotonicity regularization in AOLUT can be defined as:
[…]
where $g(\cdot)$ denotes the ReLU activation function. Similarly, $\mathcal{L}_{m}^{PG}$ and $\mathcal{L}_{m}^{AO}$ are defined in the same way as in Equation 37.
The final loss functions are as follows:
$\mathcal{L}=\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{s}+\lambda_{m}\mathcal{L}_{m},$ (11)
where the two constant parameters $\lambda_{s}$ and $\lambda_{m}$ are used to control the effects of the smoothness and monotonicity regularization terms, respectively. In our experiments, we empirically set $\lambda_{s}$ = 0.0001 and $\lambda_{m}$ = 10. More details about the loss functions can be found in Section A.5
Figure 3: Visual comparison on WorldView-III dataset. The last row visualizes the MSE residues between the pan-sharpening results and the ground truth.
4 Experiments
4.1 Datasets
Remote sensing datasets from three satellites are used in our experiments, including WorldView-II (WV2), GaoFen2 (GF2) and WorldView-III (WV3). Due to the absence of high-resolution multispectral ground truth images in these datasets, we generate the training set using the Wald protocol tool waldtool. Specifically, given the original MS image and its corresponding high-resolution PAN image, they are downsampled by a factor of $r$ to obtain image pairs of MS and PAN, with $r$ set to 4. During training, the original high-resolution MS image is treated as the ground truth, while the MS and PAN images serve as the input image pairs.
4.2 Implementation Details
We compare the proposed Pan-LUT model against several pan-sharpening methods on reduced-resolution scenes from WV2, WV3, and GF2 datasets. Specifically, we choose four traditional pan-sharpening techniques: Brovey brovey, IHS ihs, SFIM liu2000smoothing and GS GS, along with ten deep learning-based approaches: PNN pnn, PanNet pannet, MSDCNN msdcnn, Pan-GAN pangan, SFDI sfiin, UCGAN ucgan, PanFlow yang2023panflownet, PSCINN wang2024panpsci, Pan-Mamba panmamba and TA-DiffHQP tdiff. Several widely used image quality assessment metrics are employed to evaluate the performance of the algorithm, including peak signal-to-noise ratio (PSNR) psnr, structural similarity index (SSIM) ssim, spectral angle mapper (SAM) sam, relative dimensionless global error in synthesis (ERGAS) ergas, spectral distortion index($D_{\lambda}$), spatial distortion index ($D_{S}$) and the quality with no reference (QNR) qnr. The PyTorch framework is implemented in our experiment. During the training phase, we employ an ADAM optimizer with $\beta_{1}=\,\,0.9$ and $\beta_{2}=\,\,0.999$, to update the network parameters for 1000 epochs with a batch size of 1. The learning rate is initialized with $5\times 10^{-4}$. In parallel, a StepLR learning rate adjustment strategy is employed to reduce the learning rate by half after every 200 iterations. The sizes of PGLUT, SDLUT and AOLUT are set to 9, 9 and 9, respectively.
Table 2: Evaluation on the real-world full-resolution scenes from WorldView-II dataset. The best values are highlighted by red. The up or down arrow indicates higher or lower metric corresponding to better results.
Metrics Brovey IHS PNN MSDCNN GPPNN SFDI UCGAN PanFlow PSCINN Pan-Mamba TA-DiffHQP Ours
$D_{\lambda}\downarrow$ 0.1026 0.1110 0.1057 0.1063 0.0987 0.1034 0.1042 0.0966 0.0967 0.0966 0.0953  0.0571
$D_{S}\downarrow$ 0.1409 0.1556 0.1446 0.1443 0.1312 0.1305 0.1476 0.1274 0.1271 0.1272 0.1129  0.0640
[…]
Figure 4: Visual comparison on the real full-resolution scenes from the WorldView-II dataset. For a more detailed examination of the results, we zoomed-in view on specific parts of the images.
4.3 Comparison with Other Methods
Evaluation on Reduced-resolution Scene. The quantitative results across three datasets are presented in Table 1, with the best results highlighted in red. Compared to traditional methods, Pan-LUT achieves an average PSNR improvement of 5dB, 7dB, and 7dB across the three datasets, while maintaining inference speeds comparable to those of traditional methods. It is worth noting that the proposed Pan-LUT does not incorporate any network structure, yet it outperforms some DNN-based methods, such as PanNet and PNN, in terms of both performance and inference time. We also provide visual comparisons for the WV3 datasets, as shown in Figure 3.
Evaluation on Full-resolution Scene. To assess the performance and generalization capability of our method on full-resolution scenes under real-world conditions, we first trained Pan-LUT on the reduced-resolution WorldView-II data and then tested it on unseen full-resolution WorldView-II satellite datasets. The real-world dataset consists of 200 newly collected samples from the WorldView-II satellite for evaluation. The results are presented in Table 2. On reduced-resolution scenes, our method falls short of most DNN-based approaches in terms of performance. However, on full-resolution scenes, it outperforms all of them when considering the metrics of $D_{\lambda}$, $D_{S}$, and QNR. This demonstrates its strong generalization ability in real-world situations. Additionally, we provide a visual comparison against both traditional and DNN-based methods, as shown in Figure 4.
Computation Efficiency Comparison. We conduct three experiments to comprehensively evaluate the computational efficiency of all methods: (1) testing the maximum image size that each method can handle on 11GB and 24GB GPUs; (2) measuring their inference time on the CPU; and (3) evaluating their inference time on an RTX 2080 Ti GPU with $\times2K$ and $\times4K$ images. For each method, we record the average inference time on 100 images. As shown in Figure 1, even with a 24GB GPU, existing methods fail to process $\times8K$ images, while our method can handle $\times9K$ images on an 11GB GPU. In environments without GPU acceleration, most methods exhibit unsatisfactory inference speed. As shown in Tabel 1, Pan-LUT efficiently processes images at all resolutions. Compared to DNN-based methods, it achieves significantly faster inference speeds, while maintaining comparable speed to traditional methods. Our method easily meets the real-time processing requirements on GPUs, outperforming all other methods by a substantial margin. Notably, only PNN is capable of handling remote sensing satellite images at the $\times4K$ resolution, highlighting the superior efficiency of our approach.
4.4 Ablation Study
Figure 5: Ablation studies on different sizes of the PGLUT, SDLUT and AOLUT on the WorldView-III dataset.
Table 3: Memory requirements (K).
Module 3 6 9 17 33
AOLUT 0.97 31.10 236.20 5679.43 156541.57
SDLUT 0.08 1.30 6.56 83.52 1185.92
PGLUT 1.21 38.88 295.24 7099.28 195676.96
Table 4: Ablation study of PGLUT, SDLUT and AOLUT on the WorldView-III dataset.
[…]
(i) only PGLUT 25.2788 0.7742 0.1130 5.4852
(ii) only SDLUT 26.1433 0.8165 0.1123 5.4211
(iii) only AOLUT 28.9554 0.8677 0.1001 3.7661
(iv) PGLUT and SDLUT 29.1315 0.8795 0.0988 3.6785
(v) SDLUT and AOLUT 29.2754 0.9010 0.0934 3.6033
(vi) PGLUT and AOLUT 29.4875 0.8925 0.0900 3.4977
$\star$ Pan-LUT (Ours) 29.7376 0.9106 0.0815 3.3934
Size of Look-Up Tables. As shown in Figure 5, changing the LUT size does not lead to a significant drop in performance. This observation suggests that the effectiveness of our proposed method is not dependent on consuming extensive storage resources to increase the LUT size. First, we examine the effect of PGLUT size, denoted $N_{P}$. Performance improves with larger $N_{P}$, reaching an optimal point at values $N_{P}$ = 9. Beyond this (from 9 to 17), only a minor gain of 0.01 dB is observed, while the number of parameters increases substantially from 236K to 5M, indicating capacity redundancy. Therefore, we set $N_{P}$ = 9 as the default to balance performance with storage requirements. For SDLUT, denoted $N_{S}$, increasing $N_{S}$ from 3 to 9 improves performance, but values above 9 cause a slight performance drop. Similarly, enlarging the AOLUT size $N_{A}$ yields only minor gains but substantially increases parameters, especially beyond $N_{A}$ = 9. We thus set $N_{A}$ = 9 to balance performance with computational efficiency. We provide the parameter count for each LUT of different sizes in Tabel 5, which can be calculated as follows:
$\begin{gathered}Param_{PGLUT}=5N^{5},Param_{SDLUT}=N^{4},Param_{AOLUT}=4N^{5}.\end{gathered}$ (12)
Effectiveness of Each LUT. We further conduct ablation studies to verify the effectiveness of each LUT. Results are listed in Table 5. Our observations are as follows: 1) Comparing (i) with (iv) and (iii) with (v), SDLUT effectively captures fine-grained spatial details from the PAN image, thereby enhancing overall performance. 2) Comparing (ii) with (iv) and (iii) with (iv), PGLUT provides finer control over spectral transformation, resulting in improved performance. 3) Comparing (ii) with (v) and (i) with (vi), AOLUT demonstrates adaptive aggregation capabilities.
5 Conclusion
In this paper, we propose a novel learnable LUT framework, called Pan-LUT, which strikes an optimal balance between performance and computational efficiency for high-resolution remote sensing images in pan-sharpening. The proposed method makes it possible to process $15K\times 15K$ remote sensing images on a 24GB GPU and processes a $\times9K$ image in under 1 ms using one RTX 2080 Ti GPU. Extensive experiments on various satellite datasets demonstrate the effectiveness and efficiency of Pan-LUT.
Appendix A Appendix
The following contents are provided in the appendix:
  • Limitation and Border Impact.
  • More Technical Details.
  • More ablation study.
  • More comparison experiments.
  • More visualization about our experiments.
A.1 Limitation and Border Impact
Limitation. One limitation of existing methods is the lack of neural network integration. While combining LUTs with neural structures may offer greater modeling capacity, it often comes at the cost of computational efficiency. Conversely, LUTs alone require substantial storage space, making them less suitable for high-dimensional input processing. The space complexity of an n-dimensional LUT is $O(ND^{n})$, meaning its size increases exponentially with higher dimensions. This exponential growth severely limits its practical applicability, particularly when dealing with high-dimensional inputs. Lightweight network architectures must be further explored to facilitate efficient integration with LUTs. Meanwhile, more efficient LUT designs are essential for processing high-dimensional inputs.
A.2 Details of PGLUT
Input: Suppose that the MS image ($MS\in R^{H/r\times W/r\times C}$) and the PAN image ($P\in R^{H\times W\times 1}$) are given, the input of PGLUT ($I_{PG}\in R^{H\times W\times(C+1)}$) can be formulated as:
$I_{PG}=Concat(P,MS\uparrow),$ (13)
where $Concat(\cdot)$ is the concatenation operation. $MS\uparrow$ denotes the upsampled MS image. $MS\uparrow$ and $P$ have the same spatial resolution.
Output: the output of PGLUT ($O_{PG}\in R^{H\times W\times(C+1)}$) can be formulated as:
$O_{PG}=F_{PGLUT}(I_{PG}),$ (14)
where $F_{PGLUT}(\cdot)$ denotes the lookup and pentalinear interpolation based on the PGLUT.
PAN-guided indexing strategy. In a MS image, pixels from different spatial locations may have identical values (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j}$, where $i\neq j$). The LUT maps these identical inputs to the same output. The PAN-guided indexing strategy provides a more flexible indexing mechanism for the LUT. Specifically, it additionally considers the pixels at corresponding spatial positions in the PAN image (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j},pa_{i}\neq pa_{j}$, where $i\neq j$), thereby achieving finer-grained mapping. Given an input value $I_{(w,h)}=\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image, the input index to the PGLUT based on the input value can be represented as (x, y, z, s, e). PGLUT first performs a lookup operation to locate the nearest 32 adjacent elements around the input index in the PGLUT. We use $(i,j,k,m,n)$ to denote the coordinates of a defined sampling point in PGLUT, which can be calculated as follows:
$\begin{gathered}x=\frac{pa_{(w,h)}^{I}}{V_{max}}\cdot N,y=\frac{r_{(w,h)}^{I}}{V_{max}}\cdot N,z=\frac{g_{(w,h)}^{I}}{V_{max}}\cdot N,\\ s=\frac{b_{(w,h)}^{I}}{V_{max}}\cdot N,e=\frac{nir_{(w,h)}^{I}}{V_{max}}\cdot N,\\ i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,n=\lfloor e\rfloor,\end{gathered}$ (15)
where $V_{max}$ denotes the maximum value (e.g., 255, 1023 or 2047). $\lfloor\cdot\rfloor$ denotes the floor function. Then, the offset between the input precise index $(x,y,z,s,e)$ and the computed sampling point $(i,j,k,m,n)$ can be computed:
$\begin{gathered}d_{x}=x-i,d_{y}=y-j,d_{z}=z-k,\\ d_{s}=s-m,d_{e}=e-m,\\ d_{-x}=1-d_{x},d_{-y}=1-d_{y},d_{-z}=1-d_{z},\\ d_{-s}=1-d_{s},d_{-e}=1-d_{e}.\end{gathered}$ (16)
Pentalinear Interpolation. After locating 32 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$\begin{gathered}O_{PG(x,y,z,s,e)}=d_{-x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i,j,k,m,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i+1,j,k,m,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i,j+1,k,m,n)}+d_{-x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i,j,k+1,m,n)}\\ +d_{-x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i,j,k,m+1,n)}+d_{-x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i,j,k,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i+1,j+1,k,m,n)}+d_{x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i+1,j,k+1,m,n)}\\ +d_{x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i+1,j,k,m+1,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i+1,j,k,m,n+1)}\\ +d_{-x}d_{y}d_{z}d_{-s}d_{-e}O_{(i,j+1,k+1,m,n)}+d_{-x}d_{y}d_{-z}d_{s}d_{-e}O_{(i,j+1,k,m+1,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{e}O_{(i,j+1,k,m,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{-e}O_{(i,j,k+1,m+1,n)}\\ +d_{-x}d_{-y}d_{z}d_{-s}d_{e}O_{(i,j,k+1,m,n+1)}+d_{-x}d_{-y}d_{-z}d_{s}d_{e}O_{(i,j,k,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{-s}d_{-e}O_{(i+1,j+1,k+1,m,n)}+d_{x}d_{y}d_{-z}d_{s}d_{-e}O_{(i+1,j+1,k,m+1,n)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{e}O_{(i+1,j+1,k,m,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{-e}O_{(i+1,j,k+1,m+1,n)}\\ +d_{x}d_{-y}d_{z}d_{-s}d_{e}O_{(i+1,j,k+1,m,n+1)}+d_{x}d_{-y}d_{-z}d_{s}d_{e}O_{(i+1,j,k,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{-e}O_{(i,j+1,k+1,m+1,n)}+d_{-x}d_{y}d_{z}d_{-s}d_{e}O_{(i,j+1,k+1,m,n+1)}\\ +d_{-x}d_{y}d_{-z}d_{s}d_{e}O_{(i,j+1,k,m+1,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{e}O_{(i,j,k+1,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{s}d_{-e}O_{(i+1,j+1,k+1,m+1,n)}+d_{x}d_{y}d_{z}d_{-s}d_{e}O_{(i+1,j+1,k+1,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{s}d_{e}O_{(i+1,j+1,k,m+1,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{e}O_{(i+1,j,k+1,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{e}O_{(i,j+1,k+1,m+1,n+1)}+d_{x}d_{y}d_{z}d_{s}d_{e}O_{(i+1,j+1,k+1,m+1,n+1)},\end{gathered}$ (17)
where $O_{(i,j,k,m,n)}$ represents the value of the LUT at the coordinate $(i,j,k,m,n)$.
A.3 Details of SDLUT
Input: The input image ($O_{PG}\in R^{H\times W\times{(C+1)}}$) of SDLUT is the output of PGLUT.
Output: The output of SDLUT ($O_{SD}\in R^{H\times W\times{(C+1)}}$) can be formulated as:
${O_{SD}}=F_{SDLUT}(O_{PG}),\\$ (18)
where $F_{SDLUT}(\cdot)$ denotes the lookup and quadrilinear interpolation based on the SDLUT.
Look Up. Specifically, for each channel of input, SDLUT operates by iterating through pixels one at a time, treating each as the current pixel during processing. Given the input current pixel and ajacent pixels, which denoted as $p_{(w,h)}$, $p_{(w+1,h)}$, $p_{(w,h+1)}$ and $p_{(w+1,h+1)}$ respectively, the input index to the SDLUT based on the input can be represented as $(x,y,z,s)$. SDLUT first performs a lookup operation to locate the nearest 16 adjacent elements around the input index in the SDLUT. We use $(i,j,k,m)$ to denote the coordinates of a defined sampling point in PGLUT, which can be calculated as follows:
$\begin{gathered}x=\frac{p_{(w,h)}}{V_{max}}\cdot N,y=\frac{p_{(w+1,h)}}{V_{max}}\cdot N,\\ z=\frac{p_{(w,h+1)}}{V_{max}}\cdot N,s=\frac{p_{(w+1,h+1)}}{V_{max}}\cdot N,\\ i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,\end{gathered}$ (19)
where $V_{max}$ denotes the maximum pixel value. $\lfloor\cdot\rfloor$ signifies the floor function. Then, the offset between the input precise index $(x,y,z,s)$ and the computed sampling point $(i,j,k,m)$ can be computed:
$\begin{gathered}d_{x}=x-i,d_{y}=y-j,d_{z}=z-k,d_{s}=s-m,\\ d_{-x}=1-d_{x},d_{-y}=1-d_{y},\\ d_{-z}=1-d_{z},d_{-s}=1-d_{s}.\end{gathered}$ (20)
Quadrilinear Interpolation. After locating 16 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$\begin{gathered}O_{SD(x,y,z,s)}=d_{-x}d_{-y}d_{-z}d_{-s}O_{(i,j,k,m)}+d_{x}d_{-y}d_{-z}d_{-s}O_{(i+1,j,k,m)}\\ +d_{-x}d_{y}d_{-z}d_{-s}O_{(i,j+1,k,m)}+d_{-x}d_{-y}d_{z}d_{-s}O_{(i,j,k+1,m)}\\ +d_{-x}d_{-y}d_{-z}d_{s}O_{(i,j,k,m+1)}+d_{x}d_{y}d_{-z}d_{-s}O_{(i+1,j+1,k,m)}\\ +d_{x}d_{-y}d_{z}d_{-s}O_{(i+1,j,k+1,m)}+d_{x}d_{-y}d_{-z}d_{s}O_{(i+1,j,k,m+1)}\\ +d_{-x}d_{y}d_{z}d_{-s}O_{(i,j+1,k+1,m)}+d_{-x}d_{y}d_{-z}d_{s}O_{(i,j+1,k,m+1)}\\ +d_{-x}d_{-y}d_{z}d_{s}O_{(i,j,k+1,m+1)}+d_{x}d_{y}d_{z}d_{-s}O_{(i+1,j+1,k+1,m)}\\ +d_{x}d_{y}d_{-z}d_{s}O_{(i+1,j+1,k,m+1)}+d_{x}d_{-y}d_{z}d_{s}O_{(i+1,j,k+1,m+1)}\\ +d_{-x}d_{y}d_{z}d_{s}O_{(i,j+1,k+1,m+1)}+d_{x}d_{y}d_{z}d_{s}O_{(i+1,j+1,k+1,m+1)},\\ \end{gathered}$ (21)
where $O_{(i,j,k,m)}$ represents the value of the LUT at the coordinate $(i,j,k,m)$.
Rotation-indexing strategy. In general, the performance of the SDLUT can be improved when more pixels are considered. In order to exploit more area in the image, we use a Rotation-indexing strategy in the training phase. For our SDLUT, 4 rotational ensemble with 0, 90, 180, and 270 degrees covers total $3\times 3$ pixels. Each output from the 4 rotations is defined as follow:
[…]
where $F_{SDLUT}(\cdot)$ denotes the lookup and quadrilinear interpolation based on the SDLUT.
Proof Let $(w,h)$ denote the pixel located at the w-th column and h-th row of the image. We provide a theoretical proof that Rotation-indexing strategy (RiS) extends the receptive field from $2\times 2$ to $3\times 3$. Ideally, we aim to incorporate all pixels within the local $3\times 3$ region centered at $(w,h)$, corresponding to the input index set:
$\begin{gathered}G_{tgt}=\{(w-1,h-1),(w,h-1),(w+1,h-1),(w-1,h),\\ (w,h),(w+1,h),(w-1,h+1),(w,h+1),(w+1,h+1)\}.\end{gathered}$ (23)
By default, SDLUT captures the bottom-right neighborhood of each pixel, with the corresponding input index group defined as:
$G_{r0}=\{(w,h),(w+1,h),(w,h+1),(w+1,h+1)\}.$ (24)
With the proposed RiS, SDLUT is applied to three rotated versions of the input, thereby effectively capturing a broader set of pixel neighborhoods.
$\begin{gathered}G_{r90}=\{(w,h),(w-1,h),(w,h+1),(w-1,h+1)\},\\ G_{r180}=\{(w,h),(w-1,h),(w,h-1),(w-1,h-1)\},\\ G_{r270}=\{(w,h),(w+1,h),(w,h-1),(w+1,h-1)\}.\end{gathered}$ (25)
Then, we can derive the following equation:
$G_{r0}\cup G_{r90}\cup G_{r180}\cup G_{r270}=G_{tgt}.$ (26)
A.4 Details of AOLUT
Input: The input of AOLUT is the output from the SDLUT ($O_{SD}\in R^{H\times W\times{(C+1)}}$).
Output: the output of AOLUT ($O_{AO}\in R^{H\times W\times C}$) can be formulated as:
$O_{AO}=F_{AOLUT}(O_{SD}),$ (27)
where $F_{AOLUT}(\cdot)$ denotes the lookup and pentalinear interpolation based on the AOLUT.
Given an input value $I_{(w,h)}=\{O_{SD(w,h)}^{c_{1}},O_{SD(w,h)}^{c_{2}},O_{SD(w,h)}^{c_{3}},O_{SD(w,h)}^{c_{4}},O_{SD(w,h)}^{c_{5}}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image, $\{c_{1},c_{2},c_{3},c_{4},c_{5}\}$ denote the different channels of $O_{SD}$. The input index to the PGLUT based on the input value can be represented as (x, y, z, s, e). PGLUT first performs a lookup operation to locate the nearest 32 adjacent elements around the input index in the PGLUT. We use $(i,j,k,m,n)$ to denote the coordinates of a defined sampling point in PGLUT, which can be calculated as follows:
$\begin{gathered}x=\frac{O_{SD(w,h)}^{c_{1}}}{V_{max}}\cdot N,y=\frac{O_{SD(w,h)}^{c_{2}}}{V_{max}}\cdot N,z=\frac{O_{SD(w,h)}^{c_{3}}}{V_{max}}\cdot N,\\ s=\frac{O_{SD(w,h)}^{c_{4}}}{V_{max}}\cdot N,e=\frac{O_{SD(w,h)}^{c_{5}}}{V_{max}}\cdot N,\\ i=\lfloor x\rfloor,j=\lfloor y\rfloor,k=\lfloor z\rfloor,m=\lfloor s\rfloor,n=\lfloor e\rfloor,\end{gathered}$ (28)
where $V_{max}$ denotes the maximum value (e.g., 255, 1023 or 2047). $\lfloor\cdot\rfloor$ denotes the floor function. Then, the offset between the input precise index $(x,y,z,s,e)$ and the computed sampling point $(i,j,k,m,n)$ can be computed:
$\begin{gathered}d_{x}=x-i,d_{y}=y-j,d_{z}=z-k,\\ d_{s}=s-m,d_{e}=e-m,\\ d_{-x}=1-d_{x},d_{-y}=1-d_{y},d_{-z}=1-d_{z},\\ d_{-s}=1-d_{s},d_{-e}=1-d_{e}.\end{gathered}$ (29)
Pentalinear Interpolation. After locating 32 adjacent points, an appropriate interpolation technique is applied to generate the output value using the values of these sampled points:
$\begin{gathered}O_{AO(x,y,z,s,e)}=d_{-x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i,j,k,m,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{-e}O_{(i+1,j,k,m,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i,j+1,k,m,n)}+d_{-x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i,j,k+1,m,n)}\\ +d_{-x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i,j,k,m+1,n)}+d_{-x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i,j,k,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{-e}O_{(i+1,j+1,k,m,n)}+d_{x}d_{-y}d_{z}d_{-s}d_{-e}O_{(i+1,j,k+1,m,n)}\\ +d_{x}d_{-y}d_{-z}d_{s}d_{-e}O_{(i+1,j,k,m+1,n)}+d_{x}d_{-y}d_{-z}d_{-s}d_{e}O_{(i+1,j,k,m,n+1)}\\ +d_{-x}d_{y}d_{z}d_{-s}d_{-e}O_{(i,j+1,k+1,m,n)}+d_{-x}d_{y}d_{-z}d_{s}d_{-e}O_{(i,j+1,k,m+1,n)}\\ +d_{-x}d_{y}d_{-z}d_{-s}d_{e}O_{(i,j+1,k,m,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{-e}O_{(i,j,k+1,m+1,n)}\\ +d_{-x}d_{-y}d_{z}d_{-s}d_{e}O_{(i,j,k+1,m,n+1)}+d_{-x}d_{-y}d_{-z}d_{s}d_{e}O_{(i,j,k,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{-s}d_{-e}O_{(i+1,j+1,k+1,m,n)}+d_{x}d_{y}d_{-z}d_{s}d_{-e}O_{(i+1,j+1,k,m+1,n)}\\ +d_{x}d_{y}d_{-z}d_{-s}d_{e}O_{(i+1,j+1,k,m,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{-e}O_{(i+1,j,k+1,m+1,n)}\\ +d_{x}d_{-y}d_{z}d_{-s}d_{e}O_{(i+1,j,k+1,m,n+1)}+d_{x}d_{-y}d_{-z}d_{s}d_{e}O_{(i+1,j,k,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{-e}O_{(i,j+1,k+1,m+1,n)}+d_{-x}d_{y}d_{z}d_{-s}d_{e}O_{(i,j+1,k+1,m,n+1)}\\ +d_{-x}d_{y}d_{-z}d_{s}d_{e}O_{(i,j+1,k,m+1,n+1)}+d_{-x}d_{-y}d_{z}d_{s}d_{e}O_{(i,j,k+1,m+1,n+1)}\\ +d_{x}d_{y}d_{z}d_{s}d_{-e}O_{(i+1,j+1,k+1,m+1,n)}+d_{x}d_{y}d_{z}d_{-s}d_{e}O_{(i+1,j+1,k+1,m,n+1)}\\ +d_{x}d_{y}d_{-z}d_{s}d_{e}O_{(i+1,j+1,k,m+1,n+1)}+d_{x}d_{-y}d_{z}d_{s}d_{e}O_{(i+1,j,k+1,m+1,n+1)}\\ +d_{-x}d_{y}d_{z}d_{s}d_{e}O_{(i,j+1,k+1,m+1,n+1)}+d_{x}d_{y}d_{z}d_{s}d_{e}O_{(i+1,j+1,k+1,m+1,n+1)},\end{gathered}$ (30)
where $O_{(i,j,k,m,n)}$ represents the value of the LUT at the coordinate $(i,j,k,m,n)$.
A.5 Details of Loss Function
To achieve satisfying pan-sharpening results, we propose a joint loss for network training. Suppose the batch size is $T$. We first utilize the $\mathcal{L}_{1}$ loss:
$\mathcal{L}_{mse}=\frac{1}{T}\sum_{t=1}^{T}\|HRMS_{t}-GT_{t}\|^{2},$ (31)
where $HRMS$ and $GT$ denote the network output and the corresponding ground truth, respectively.
To enhance the stability and robustness of the learned LUTs, we incorporate smoothness regularization $\mathcal{L}_{s}$ and monotonicity regularization $\mathcal{L}_{m}$:
$\mathcal{L}_{s}=\mathcal{L}_{s}^{PG}+\mathcal{L}_{s}^{SD}+\mathcal{L}_{s}^{AO},$ (32)
$\mathcal{L}_{m}=\mathcal{L}_{m}^{PG}+\mathcal{L}_{m}^{SD}+\mathcal{L}_{m}^{AO},$ (33)
where $\mathcal{L}_{s}^{PG}$, $\mathcal{L}_{s}^{SD}$, and $\mathcal{L}_{s}^{AO}$ denote the smoothness regularizations for PGLUT, SDLUT, and AOLUT, while $\mathcal{L}_{m}^{PG}$, $\mathcal{L}_{m}^{SD}$, and $\mathcal{L}_{m}^{AO}$ represent the monotonicity regularizations for PGLUT, SDLUT, and AOLUT, respectively.
The Smoothness Regularization of PGLUT and AOLUT:
$\begin{gathered}\mathcal{L}_{s}^{PG},\mathcal{L}_{s}^{AO}=\sum_{O\in\{l,o,c,a,e\}}\sum_{i,j,k,m,n=0}^{N-1}(\left\|O_{(i+1,j,k,m,n)}\right.\\ -O_{(i,j,k,m,n)}\Big\|^{2}+\left\|O_{(i,j+1,k,m,n)}-O_{(i,j,k,m,n)}\right\|^{2}\\ +\left\|O_{(i,j,k+1,m,n)}-O_{(i,j,k,m,n)}\right\|^{2}\\ +\left\|O_{(i,j,k,m+1,n)}-O_{(i,j,k,m,n)}\right\|^{2}\\ +\left\|O_{(i,j,k,m,n+1)}-O_{(i,j,k,m,n)}\right\|^{2}).\end{gathered}$ (34)
The Monotonicity Regularization of PGLUT and AOLUT:
$\begin{gathered}\mathcal{L}_{m}^{PG},\mathcal{L}_{m}^{AO}=\sum_{O\in\{l,o,c,a,e\}}\sum_{i,j,k,m,n=0}^{N-1}[g(O_{(i,j,k,m,n)}\\ -O_{(i+1,j,k,m,n)})+g(O_{(i,j,k,m,n)}-O_{(i,j+1,k,m,n)})\\ +g(O_{(i,j,k,m,n)}-O_{(i,j,k+1,m,n)})\\ +g(O_{(i,j,k,m,n)}-O_{(i,j,k,m+1,n)})\\ +g(O_{(i,j,k,m,n)}-O_{(i,j,k,m,n+1)})],\end{gathered}$ (35)
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m,n)}$ is the corresponding output for the defined sampling point $(i,j,k,m,n)$ in LUT. $g(\cdot)$ denotes the ReLU activation function.
The Smoothness Regularization of SDLUT:
[…]
The Monotonicity Regularization of SDLUT:
[…]
where $N$ represents the number of bins in each dimension of the LUT. $O_{(i,j,k,m)}$ is the corresponding output for the defined sampling point $(i,j,k,m)$ in LUT. $g(\cdot)$ denotes the ReLU activation function.
The final loss function is as follows:
$\mathcal{L}=\mathcal{L}_{1}+\lambda_{s}\mathcal{L}_{s}+\lambda_{m}\mathcal{L}_{m},$ (38)
where the two constant parameters $\lambda_{s}$ and $\lambda_{m}$ are used to control the effects of the smoothness and monotonicity regularization terms, respectively. In our experiments, we empirically set $\lambda_{s}$ = 0.0001 and $\lambda_{m}$ = 10.
A.6 Selection of regularization parameters.
As shown in Table 6,  6, we vary the $\lambda_{s}$ and $\lambda_{m}$ to determine the optimal parameters ($\lambda_{m}=10$ and $\lambda_{s}=0.0001$). A large $\lambda_{s}$ (e.g., >0.0001>0.0001) results in worse PSNR, as the smooth regularization limits the flexibility of LUT transformations. In contrast, the PSNR is insensitive to the choice of $\lambda_{m}$ since monotonicity is a natural constraint to LUTs.
A.7 More Comparison With Traditional methods
We further compare our method with additional traditional approaches, as shown in Table 7. While significantly surpassing traditional methods in performance across the three datasets, our approach maintains a comparable processing speed, highlighting its practical efficiency. Notably, some traditional methods are not only time-consuming but also inefficient.
A.8 More Visualization about Experiments
As shown in Figure 6 and Figure 7, extensive visual comparisons are provided for the WV2 and GF2 datasets. More visual comparisons are provided in the supplementary material.
Table 5: Ablation study on $\lambda_{m}$
Metrics 0 0.1 1 10 100
PSNR 29.5321 29.5421 29.6241 29.7376 29.4112
Table 6: Ablation study on $\lambda_{s}$
Metrics 0 0.00001 0.0001 0.001 0.01
PSNR 29.4746 29.4889 29.7376 29.5423 29.5213
Table 7: Quantitative comparison across three satellite datasets with traditional methods. The best outcomes are highlighted in red. $\uparrow$ indicates better performance with increasing values, while $\downarrow$ signifies improved performance with decreasing values.
WorldView-II GaoFen2 Worldview-III Inference (ms)
[…]
PCA 20.3542 0.7002 0.3741 10.8524 19.2933 0.7974 0.3908 14.3228 20.4455 0.5263 0.2693 10.3129 0.21 0.23
Brovey 35.8646 0.9216 0.0403 1.8238 37.7974 0.9026 0.0218 1.3720 22.5060 0.5466 0.1159 8.2331 0.28 0.33
IHS 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616 0.23 0.26
SFIM 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730 0.32 0.47
Wavelet 34.9827 0.8806 0.0481 2.0907 35.7502 0.8213 0.0283 2.0148 21.8551 0.5216 0.1368 9.1158 13.37 21.73
GS 35.6376 0.9176 0.0423 1.8774 37.2260 0.9034 0.0309 1.6736 22.5608 0.5470 0.1217 8.2433 0.75 0.87
GSA 35.3574 0.9219 0.097 1.7401 35.948 0.8779 0.0368 1.9257 21.8845 0.5458 0.1394 9.0781 0.73 0.76
GFPCA 34.5580 0.9038 0.0488 2.1400 37.9443 0.9204 0.0314 1.5604 22.3344 0.4826 0.1294 8.3964 0.42 0.66
PRACS 34.9671 0.9063 0.0414 1.8725 36.2015 0.8902 0.0372 1.8312 22.4452 0.5535 0.1373 8.2961 2.44 4.75
AWLP 32.2402 0.8709 0.0457 2.4077 37.2183 0.8917 0.0281 1.5966 21.5792 0.5323 0.1260 9.0636 8.39 14.17
MTF-GLP-HPM 31.3946 0.8722 0.0492 3.3040 37.9443 0.9204 0.0314 1.5604 21.1033 0.5505 0.1233 9.8406 4.22 6.09
Pan-LUT (Ours) 40.8555 0.9633 0.0254 1.0339 43.7466 0.9726 0.0169 0.8027 29.7376 0.9106 0.0815 3.3934 0.38 0.54
Figure 6: Visual comparison on WV2 dataset. The last row visualizes the MSE residues between the pan-sharpening results and the ground truth.
Figure 7: Visual comparison on GaoFen2 dataset. The last row visualizes the MSE residues between the pan-sharpening results and the ground truth.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction sections offer a comprehensive discussion of the manuscript’s context, intuition, and ambitions, as well as its contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the work are discussed by authors in the Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: For each theoretical result, the paper provides the full set of assumptions and a complete (and correct) proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The pipeline of the methods and the details of experiments are presented with corresponding reproducible credentials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All utilized data are sourced from open-access platforms. The code, which will be made publicly available, is uploaded as a zip file.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so "No" is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The pipeline of the methods and the details of experiments are presented with corresponding reproducible credentials.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The results contain the standard deviation of the results over several random runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The details of experiments are presented with corresponding reproducible credentials.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original owners of assets, including data and models, used in the paper, are properly credited and are the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The new assets introduced in the paper are well documented and provided alongside the assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The core method development in this research dose not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.